OpenStudy (anonymous):

what is (1/secx+1)+(1/secx-1)=??

OpenStudy (anonymous):

secx= 1/cosx then it is obvious that 1/secx=cosx so we have (cosx+1)(cosx-1) which is equal with (cosx^2-1).

OpenStudy (turingtest):

actually the terms are added in the post, so\[({1\over\sec x}+1)+({1\over\sec x}-1)=(\cos x+1)+(\cos x-1)=2\cos x\]

OpenStudy (anonymous):

oh sorry, didn`t see that they r added. so yeah turningtest is right.